The local community college theater put on a production of chicago. There were tickets sold, some for $16(non student price) and others for $12(student price). If the receipts for one performance totaled $2640, How much of each type of ticket were sold?


1
Expert's answer
2021-10-06T10:08:36-0400

Let "x=" the number of tickets sold for non students, "y=" the number of tickets sold for students.

The receipts for one performance totaled $2640


"16x+12y=2640"

"4x+3y=660"

If there were 186 tickets sold


"x+y=186=>y=186-x"

Substitute


"4x+3(186-x)=660"

"x=102"

"y=84"

There were 102 tickets for non students and 84 tickets for students sold.